the sum of three consecutive odd numbers is 63.Find them.

Answer by JulietG(1812)

B=A+2
C=A+4
where A is the smallest number (they're consecutive odd so they'll be 2 apart)
A + B + C = 63
Substitute the known values of B & C from the top equations into the bottom
A + B + C = 63
A + (A+2) + (A+4) = 63
3A + 6 = 63
Subtract 6 from each side
3A = 57
Divide each side by 3
A = 19
If A is 19, then B is 21, and C is 23
19+21+23=63


There's a faster way.
Take 63 and divide it by the number of values
63/3=21
That will be your middle number, with a number on each side of it.